OpenStudy (abhisar):

Electronegativity and Bond polatity @superhelp101

OpenStudy (abhisar):

\(\begin{array}{|c|c|c|c|c|c|c|c|} \hline \rm F& \rm O& \rm Cl& \rm N& \rm Br& \rm C& \rm S& \rm H& \rm Na& \rm K\\ \hline 4.0&3.5&3.0&3.0&2.8&2.5&2.5&2.1&0.9&0.8\\ \hline\end{array}\)

OpenStudy (abhisar):

\(\bullet\) If the electronegativity difference (usually called ΔEN) is less than 0.5, then the bond is nonpolar covalent. \(\bullet\) If the ΔEN is between 0.5 and 1.6, the bond is considered polar covalent If the ΔEN is greater than 2.0, then the bond is ionic. \(\bullet\) If the ΔEN is between 1.6 and 2.0 and if a metal is involved, then the bond is considered ionic. If only nonmetals are involved, the bond is considered polar covalent.

OpenStudy (abhisar):

Let's see a few examples \(\Large\text{Methane (CH\(_4\))}\) In C-H bond ΔEN (Electronegativity difference) = 2.5-2.1=0.4. Now since it is < 0.5 it will be non-poar (Rule1). \(\color{blue}{\text{Thus valence electrons are shared equally between C & H}}\)

OpenStudy (abhisar):

\(\Large\text{Water(H\(_2\)O)}\) In O-H bond ΔEN = 3.5-2.1=1.4. Now since 0.5<1.4<1.6 it will be a polar bond (Rule2). \(\color{blue}{\text{Thus valence electrons are not equally shared and spend more time near} \\ \text{oxygen atom because it has higher electrronegativity value.}}\)

OpenStudy (abhisar):

\(\Large\text{Hydrogen Fluoride(HF) & Sodium Bromide(NaBr)}\) In H-F bond ΔEN = 3-2.1=0.9. In Na_Br bond ΔEN=0.9-2.8=1.9. Now since in both ΔEN = 1.9>1.6<2.0. Rule 4 will be applied. Since both H & F are non-metals HF will be Polar covalent. Na is metal so NaBr will be ionic.

OpenStudy (superhelp101):

oh sorry. Thank you

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (eric_d):

This explanation is helpful to me.... @Abhisar

OpenStudy (abhisar):

That's great to hear \(\color{green}{\huge\ddot\smile}\)

OpenStudy (anonymous):

Great work @abhisar :)

OpenStudy (abb0t):

Is this a question?